A Complete Bibliography of Publications in Structured Programming and Software — Concepts and Tools

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

06 January 2024 Version 2.00

Title word cross-reference

2 [Bla89, MW91b, MW91a, Ode89].

3 [Dei98, Gut94].

80 [ETJJ93].

Abstract [CIdLS93, KT90, SGHR93]. Abstraction [Mos95, Sch91]. Action [BS91, BS96]. Active [PPSS95]. Ada [Dob92b]. Adaptive [Dei97]. Adding [YF95a]. aided [WFC00]. Algorithm [CM95, Fuc93]. Algorithms [Bro90, CP97, Har90, Pir92, Sal97, Sch92]. Alightweight [DP98]. Alternative [Wir96]. Amalgamating [CM92]. Analysis [CL97, CS97, ETJJ93, SMC96, KK92]. Analyzer [AHP96]. Animator [CM95]. Anon [Coo98]. Anon-line [Coo98]. Application [Dei97, WGM89]. Applications [MGS95, McC97]. Applied [BC90b, Eji90, WP94]. Applying [ETJJ93]. Approach [CT94, CM92, CLZ97, EBP93, KSL^{+}95, Ode89, Ros92, TR95, Yea89, HNR00, Lee00]. Approaches [BKS94]. Apsects [SGHR93]. architectural [PS98]. Architecture [Coo96, GP95a]. Article [Bro92]. Aspects [Mil91]. atomic [San00]. Authors [Bro92, Pom00]. await [San00].

Balanced [Bro90]. BaLinda [YF95a]. Based [CT94, Cha91, CFGGR91, Dei97, Pre97, TR95, BMM92, HNR00, Lee00].

Behaviour [Ros92]. between [MW91b]. Beyond [Eji90]. Binary [XD97, Yea89]. BLAS [SUL93]. Breaking [Low96]. Building [Riz93, Spi93]. Business [Dei98, KSL^{+}95, Gol98].

C [Aba89, CL97, Coo98, DP98, ETJJ93, Sak92, SUL93, SD91]. C/UNIX [SD91]. Case [CNSL96, Fra93, GJ91, SUL93]. Challenge [GB93]. Changeable [Riz93]. characterizes [BDS98]. Charts [AHP96]. Checking [Vol95]. Chiron [Sch90]. choices [TW98]. Class [SUL93, TR95, Coo98, DP98]. Class-Based [TR95]. Clean [KT90]. Clients [Cre95]. Closed [Hre97]. Closed-Loop [Hre97]. Code [Aba89, Gup90, Zim95]. Coding [Dei97]. Coherent [Vol95]. Collection [Har90]. COM [PS98]. COM/DCOM [PS98]. Comfortable [Gro90]. commands [San00]. Comparison [BPS89, Bös93, Har90, Szy98]. Compile [Aba89]. Compiled [Zim95]. Compiler [GJ91, Gup90]. Completeness [BK94]. Complexity [SS92]. Component [Bro97, CFGGR91, Pre97, Bro98, BDS98, HNR00, Szy98]. Component-Based [Pre97, HNR00]. components [PS98]. Compression [Dei97]. Computation [BK94, YF95b]. Computer [WFC00]. Computer-aided [WFC00]. Computers [Dob92b, KSA94]. Computing [Tur92]. Concept [Bro97, Bro98]. Concepts [Bla91, PBK^{+}91, Rei94, WH94]. Concurrent [CRE94, Ros92, San00]. Configuration [NB96]. Consecutive [Mos95]. Construct [GV90]. Construction [MGS95, McC97]. Constructs [GK92]. Control [Hre97, Rei94, ZWC98]. Controlled [NB96]. Conventional [BC90b]. Conversations [MG93]. Conversion [XD97]. CORBA [PS98]. Coroutines [Pad90]. Coupling [SUL93]. Course [GB93, Bös98]. Critical [Mil91].

Crystal [ETJJ93]. Cycle [BK89].

Darker [Sak92]. Data

[CL97, CIdLS93, Dei97, Gro91b, KT90]. Database [Coo96, SGHR93]. DCOM [PS98]. Dear [Pom00]. Debugging [Gup90]. Decimal [XD97]. Decomposition [Fuc93]. definable [SI98]. Definition [CC97]. Depot4 [Lam98]. Derandomization [Sal97]. Derivation [XD97]. Deriving [Fuc93, RBP89]. Descent [Gro90]. describing [Mos95]. Description [CFGGR91, Yea89]. Design [BC90b, Gro91a, Mey89, Ode89, RBP89, Sch91, Sta96, Tem91, WGM89, TW98]. Designing [Wir89]. Detection [Bre96, CL97, Hol96]. develop [San00]. Development [Gut93, PBK^{+}91, PPSS95, Pre97, RM90, HNR00, TW98, WFC00]. Diagram [KSA94]. Differences [MW91b]. Direct [MGS95]. Display [Pri89]. Display-oriented [Pri89]. Distributed [CP97, CT94, CNSL96, LS95, MGS95, McC97, SMC96, PS98]. Document [WY94]. Documents [BVW97, WH94]. Domesticating [GK92]. Dynamic [CL97, CLM96, PC98]. dynamically [Lam98]. Dynamics [WP94].

Early [Hol96]. Editor [KSA94]. Editorial [Pom89, Pom90, PR94, Pom98]. Editors [WY94]. Efficient [Gro90]. EGS1 [GP95a]. Eiffel [Mey89]. embedding [Rad98]. Embeddings [RBP89]. Emerging [Szy98]. Empirical [GB93]. Enactable [CC97]. End [MGS95, McC97]. End-User [MGS95, McC97]. Engineering [GB93, MS95, Pre97, Lee00]. Enhancing [BK89]. Environment [BG92, BC90a, CP97, KSL^{+}95, SD91]. Environments [BKS94]. Error [CL97, Gro90]. EUPHORIA [MGS95]. Evaluation [GJ91, Mil91, TW98]. Examples [Mor90]. Exception [EBP93, Pad90]. Executable [GP95a, Fuc92]. Exercise [Spi93]. Existing [SUL93]. experience [Coo98]. Experiences [Dob92b]. Explore [Sch92]. Expression [Mor90]. extended [DP98]. Extending [Cre95, Tur92]. Extensible [Coo96, Fra95, Lam98]. Extension [Fra95].

Family [Sch92]. Fault [Hol96]. favorite [Bös98]. FDR [Low96]. Feature [Bre96]. fighting [ZWC98]. Files [Fra93]. Find [RM90]. fire [ZWC98]. fire-fighting [ZWC98]. First [Lin89a, Bös98]. first-course [Bös98]. Fixing [Low96]. Flow [CL97]. Flowgraph [Fuc93]. Foreword [MS96]. Formal [GP95b]. Formalization [SMCB96]. Formally [CFGGR91]. Foundation [CFGGR91]. Framework [Dei98, PPSS95, WGM89]. Francisco [Hen98]. Fresh [Fra97]. Functional [GK92]. Future [Gut94].

GADGETS [Mar91]. Garbage [Har90]. Generation [GJ91, Tre96, MKVM00]. generator [Lam98]. Generic [Bro90]. Genetic [Sal97]. Graphic [KSA94, Mil91]. Graphical [GP95a]. Graphics [CM92]. Guidance [PPSS95].

Handling [EBP93, Pad90]. Haskell [Spi93]. Height [Bro90]. Heterogeneous [SMC96]. History [Yea89]. History-sensitive [Yea89]. Hoare [Fuc92]. Homogeneous [Sch90]. Hyperdocuments [Dob96].

IBM [Hen98]. Imperative [Epp93, GK92]. Implementation [Bla89, CLM96, EBP93, SGHR93, Spi93, Tem91, WGM89, Coo98]. Implementing [Spi93]. Improved [Fuc93]. Improving [Sal97]. Incremental [Chr89, SMCB96]. Indexes [Knu94]. Industrial [SMCB96]. Inference [HSA95]. Inheritance [BB95, Sch92]. Initialization [BS00]. Input [Ode89]. Input/Output [Ode89]. Inputs [Tre96]. Integrating [Dob92a, Sta96, SD91]. Integration [WY94]. Intelligent [Bre96]. Interactions [Bre96]. Interactive [BC90a]. Interaface [Sta96]. Interface [KT90, Mar91, SGHR93, Heu98]. Interfaces [Cre95, Gut89, MGS95]. Intermediate [Bro92, HP91]. Introducing [GB93]. Introduction [CP96, Eji90, Lin89b]. Invalidating [Cre95]. Investigation [GB93]. Issues [Gro91a]. Iterative [YF95b]. Iterators [Kof93].

Java [BS00, Bös98, PS98, San00]. JSD [Ros92].

Key [Low96, SMCB96]. Knowledge [NB96]. Knowledge-Controlled [NB96]. Knuth [XD97].

Lagoona [Fra97]. Language [Cha91, CLM96, CRE94, CFGGR91, CC97, Dob96, Fra97, GM94, Goo97, GV90, Gro91a, Gut93, HI94, LP92, Mil91, MS90, MT89, MW91a, Sch90, Spi93, War94, Bös98, Dob92a]. Language-Oriented [War94]. Languages [BPS89, Epp93, Rec90, YF95a]. Large [Fra95]. Length [RM90]. Letter [Bro92]. Levels [Mos95]. Library [Ode89, DP98]. Life [BK89]. Line [BVW97, Coo98]. Linear [Bri90, DF89]. LIPED [BG92]. LIST [Epp93]. Literate [BG92, BC90a, BC90b, CS97, Knu94, Lin89a, Lin89b, Tun89]. Local [HP91]. Logic [Chr93, CLM96, GM94, Fuc92]. Longest [RM90]. Look [Fra97, Lin89a]. Loop [Hre97]. LZ [Dei97]. LZ-Based [Dei97].

M [RM90]. M-Sequence [RM90]. Macrotec [KSL^{+}95]. Management [Mar91, NB96, DP98]. Manipulation [MGS95, SI98]. Matching [Pir92]. Math [BVW97]. Math/pad [BVW97]. Mathematical [BVW97, Bro97, Rév91, Bro98]. Measuring [SS92]. Mechanism [EBP93]. Message [AHP96]. Metamodels [Mos95]. Method [Tun89]. Methodological [CT94]. Methodology [BB95, Chr93]. Methods [GP95b]. Metrics [Eji90]. Mini [Knu94]. Mini-Indexes [Knu94]. minimal [Heu98]. MINOS [Ode89]. Model [BK94, Lee00, Gol98]. Model-based [Lee00]. Modeling [CM92, CNSL96]. Modelling [Ros92, WP94]. Models [Rec90]. Modest [MT89]. Modula [Bös93, EBP93, KT90, Pad90, Bla89, Ode89]. Modula-2 [EBP93, KT90, Pad90, Bla89, Ode89]. Modula-3 [Bös93]. Modular [BS96, CLM96]. Module [Cre95, Gut89, Rad98]. Modules [Lam97]. Multimedia [McC97, PC98]. Multiprocessing [Wir96]. my [Bös98].

Needham [Low96]. Needham-Schroeder [Low96]. Nested [Pad90]. network [ZWC98]. Networks [Bre96]. Neural [ZWC98]. Notation [Pri89, Rév91]. Numeric [CRE94].

Oberon [Bös93, Hre97, LS95, Lam97, CM92, Gut94, MT89, MW91b, MW91a, Tem91]. Oberon-2 [Bös93, MW91b, MW91a]. Object [BKS94, BPS89, Bla91, Bol93, CT94, Cha91, CLZ97, CC97, Fra97, GB93, Gro91a, LS95, Mey89, Mos95, MT89, Riz93, Sch90, SGHR93, SD91, WGM89, DP98, MT89]. Object-Based [CT94, Cha91]. Object-Orientation [Fra97]. Object-Oriented [BKS94, BPS89, Bla91, Bol93, CLZ97, CC97, GB93, Mey89, MT89, Riz93, Sch90, SGHR93, WGM89]. Objects [Bla89, LS95, YF95a, PS98]. Observations [Cel89]. Occam [LP92]. Omega [Bla91]. On-Line [BVW97]. Opaque [CP96]. Open [GP95b]. Optimized [CLM96]. Ordinary [Rév91]. Orientation [Fra97]. Oriented [BKS94, BPS89, Bla91, Bol93, CLZ97, CC97, GB93, Gro91a, Mey89, Mos95, MT89, PBK^{+}91, Riz93, Sch90, SGHR93, SD91, War94, WGM89, Dob92a, Pri89]. Output [Ode89]. Outputs [Tre96]. Overloading [Vol95].

pad [BVW97]. Paradigm

[GB93, Pre97, Tur92, Wir96]. Parallaxis [Brä89]. Parallel [Bol93, CLZ97, MS90, YF95a, KK92]. Parameterizing [Aba89]. Parameters [HP91]. Parsers [Gro90]. parsing [KK92]. Partial [GJ91]. Pascal [HSA95]. Pattern [Pir92]. Pattern-Matching [Pir92]. PBOOL [CC97]. Pen [KSA94]. Performance [Sal97]. persistence [Dob92a]. Persistent [LS95, DP98]. Personal [Dob92b]. Perspective [HI94]. pipeline [KK92]. Plus [YF95a]. Portable [MS90]. Practices [Cel89]. Preparation [BVW97]. principled [Lee00]. Principles [Eji90]. Priorities [CNSL96]. problems [BS00]. Procedures [HP91]. Processes [CC97]. Processing [YF95b]. Program [RM90, RBP89, XD97, SI98, WFC00]. Programming [BKS94, BG92, BPS89, Bla91, Bol93, Brä89, BC90a, BC90b, CT94, Cel89, Cha91, Chr93, CLZ97, Eji90, Fra97, GM94, Gro91a, Gut93, Lin89a, Lin89b, Mey89, MS90, Mor90, MW91a, Rec90, SD91, Tun89, War94]. Programs [CS97, Chr93, Knu94, TR95, Fuc92, San00, SS92]. Projects [WP94]. Prolog [GM94]. Proof [XD97]. Proposition [Mos95]. Protocol [Fra95, Low96, PC98]. Prototypes [Bla91]. Prototyping [BK89, PBK^{+}91, BMM92, Dob92a]. Prototyping-Oriented [PBK^{+}91]. Public [Low96]. Public-Key [Low96].

QPC [Bol93]. Querying [MS95].

Quiescence [Tre96]. random [MKVM00]. rapid [BMM92]. readers [Pom00]. Real [Cha91, WFC00]. Real-Time [Cha91, WFC00]. reasoning [Lee00]. Recovery [Gro90]. Recursive [Gro90, RBP89]. Rediscovered [Bri90]. Reengineering [KSL^{+}95]. Refinement [BS91, Mor90, Heu98]. Refining [Ros92]. Relational [SGHR93]. Relations [Yea89]. Reorganized [Gup90]. Repetitive [Tre96]. Representations [Sta96]. Requirements [BKS94]. retrieval [BMM92]. Reusability [LP92]. Reusable [Riz93, Sch91]. Reuse [CS97, CFGGR91, CC97, Gol98]. Reverse [MS95]. Revisited [DF89, Sak92]. Rigorous [Ros92]. RMI [PS98]. Road [Mey89]. robot [ZWC98]. Robust [Kof93]. Role [Gut89]. Routines [MG93]. Rule [SD91]. Rule-Oriented [SD91]. Run [Aba89]. RVGEN [MKVM00].

Safe [Bla91]. San [Hen98]. SAPR [Dei98]. SAPR/3 [Dei98]. Sather [Goo97]. Sather-K [Goo97]. Scheduling [CP97, Gup90]. Scheme [Dei97]. Schroeder [Low96]. Scratch [Wir89]. Seamless [WGM89]. Search [Bri90, DF89]. Selection [NB96]. Semantic [BMM92, WY94, KK92]. Semantic-based [BMM92]. sensitive [Yea89]. Sequence [AHP96, Mor90, RM90]. set [Dob92a]. set-oriented [Dob92a]. Seymour [MS90]. Sharing [Chr89]. Side [Sak92]. SIMD [Brä89]. Simple [Coo96, Heu98]. Simulating [CP97]. simultaneous [SI98]. Smalltalk [ETJJ93]. Smalltalk-80 [ETJJ93]. Software [BK89, BC90b, CP97, Eji90, Fra95, GP95a, GB93, Gut94, NB96, PBK^{+}91, Pre97, Rei94, Riz93, SUL93, WP94, WH94, BMM92, BDS98, HNR00, Lee00, Szy98, TW98]. Software-Engineering [GB93]. Software-Systems [Fra95]. Some [Dob92b]. SPARC [Tem91]. SPARC-Oberon [Tem91]. Specification [GP95a]. Specifications [Chr93, Ros92, RBP89, Fuc92]. Specifying [Dob96]. Speculative [YF95b]. Stability [SS92]. stacks [PC98]. Standard [Epp93]. StarLite [CM95]. State [TR95]. State-Based [TR95]. Statistics [Lam97]. Stepwise [BS91]. strategic [Szy98]. String [RM90]. Structure [Chr89]. Structured [Brä89, CLZ97, Eji90, GM94, Mey89, Riz93, Tun89]. Structures [Gro91b]. Structuring [Fra95]. Student [Cel89]. Studies [HI94]. Study [CNSL96, GJ91, McC97, SUL93]. Stuudyy [Pir92]. subscribers [Pom00]. Substitution [Zim95]. Success [SMCB96]. Support [CRE94, Gro91b, KSL^{+}95, BMM92]. Symbol [Fra93, Sch91]. Symbolic [CRE94]. Symmetric [Bro90]. Symmetry [ETJJ93]. Synapse [Cha91]. Syntax [Pri89]. Synthesis [Chr93]. System [BVW97, Coo96, Gut94, Lam97, Mar91, SD91, WP94, Wir89, SI98]. Systematic [TW98]. Systems [BS91, BS96, Chr89, CNSL96, CRE94, Fra95, GP95b, LS95, Mos95, SMC96].

Table [Sch91]. Task [CP97].

Task-Scheduling [CP97]. Tasks [Wir96]. Technique [Fra95]. Techniques [Rei94]. technologies [Szy98]. Technology [Gut94]. Test [Tre96]. Testing [TR95]. Thought [Rec90]. Threads [Wir96]. Three [Har90]. Time [Aba89, Cha91, WFC00]. Tool [CT94, GP95a, Gro91b, Spi93, Heu98, MKVM00]. Tools [Hol96, PBK^{+}91, WY94, WH94, TW98]. TOROS [SD91]. Trace [Gup90]. Tragedy [Gut93]. Transform [GV90]. Translating [Rév91]. translators [Lam98]. Transmission [LS95]. Tree [Fuc93]. Trees [Bro90]. Turbo [HSA95]. Type [Bla91, Coo96, Epp93, HSA95, Vol95, Zim95]. Type-Safe [Bla91]. Types [CP96, KT90].

Understandable [Riz93]. uniform [Bro98].

Universal [Fra93]. UNIX [SD91]. Usability [McC97]. Use [Bro97]. User [Mar91, MGS95, McC97, Sta96, Heu98, SI98]. user-definable [SI98]. Using [CM92, Chr93, CNSL96, LS95, Low96, San00, Sch92].

Variable [Bro92]. Variables [HP91]. variates [MKVM00]. Variations [Gut89]. Verification [Bre96, SMC96]. Verifying [CNSL96]. Version [NB96, Rei94]. versus [Wir96]. view [PS98]. Viewpoint [CS97]. Views [CIdLS93, SI98]. VinEd [SI98]. Vision [Gut94]. Visual [BK94, HI94, HNR00]. Visualization [CM92]. Visualizing [MS95].

WEB [SS92]. Welcome [Pom89]. without [Cre95]. Workflow [Sta96]. World [GK92].

References

Abacus:1989:PCC

[Aba89] Alex Abacus. Parameterizing C code at compile and run time. Structured Programming, 10(4):209–214, 1989. CODEN STPGEM. ISSN 0935-1183.

Alur:1996:AMS

[AHP96] Rajeev Alur, Gerard J. Holzmann, and Doron A. Peled. An analyzer for message sequence charts. Software — Concepts and Tools, 17(2):70–77, 1996. CO-DEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Breu:1995:MI

[BB95] Ruth Breu and Michael Breu.

A methodology of inheritance. Software — Concepts and Tools, 16(3):113–123, 1995. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Brown:1990:IEL

[BC90a] Marcus E. Brown and Bart Childs. An interactive environment for literate programming. Structured Programming, 11(1): 11–26, 1990. CODEN STPGEM. ISSN 0935-1183.

Brown:1990:LPA

[BC90b] Marcus E. Brown and David Cordes. Literate programming applied to conventional software design. Structured Programming, 11(2):85–98, 1990. CO-DEN STPGEM. ISSN 0935-1183.

Broy:1998:WCS

[BDS98] Manfred Broy, Anton Deimel, and Clemens Szyperski. What characterizes a (software) component? Software — Concepts and Tools, 19(1):49–56, March 1998. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic). URL https:/

/link.springer.com/article/
10.1007/s003780050007.

Bishop:1992:LPL

[BG92] Judy M. Bishop and Kevin M. Gregson. Literate programming and the LIPED environment. Structured Programming, 13(1): 23–34, 1992. CODEN STPGEM. ISSN 0935-1183.

Bischofberger:1989:ESL

[BK89] Walter R. Bischofberger and Rudolf K. Keller. Enhancing the software life cycle by prototyping. Structured Programming, 10(1):47–59, 1989. CO-DEN STPGEM. ISSN 0935-1183.

Brown:1994:CVC

[BK94] Timothy B. Brown and Takayuki Dan Kimura. Completeness of a visual computation model. Software — Concepts and Tools, 15(1):34–48, 1994. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Bischofberger:1994:OOP

[BKS94] Walter R. Bischofberger, Thomas Kofler, and Bruno Schäffer. Object-oriented programming environments: Requirements and approaches. Software — Concepts and Tools, 15(2):49–60, 1994. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Blaschek:1989:IOM

[Bla89] Günther Blaschek. Implementation of objects in Modula-2. Structured Programming, 10 (3):147–155, 1989. CODEN STPGEM. ISSN 0935-1183.

Blaschek:1991:TSO

[Bla91]  Günther Blaschek.  Type-

safe object-oriented programming with prototypes — the concepts of Omega. Structured Programming, 12(4):217–226, 1991. CODEN STPGEM. ISSN 0935-

Boudriga:1992:SBS

[BMM92] Noureddine Boudriga, Ali Mili, and Roland Mittermeir. Semanticbased software retrieval to support rapid prototyping. Structured Programming, 13(3):109– 127, 1992. CODEN STPGEM. ISSN 0935-1183.

Boles:1993:POO

[Bol93] Dietrich Boles. Parallel objectoriented programming with QPC++. Structured Programming, 14(4):158–172, 1993. CO-DEN STPGEM. ISSN 0935-1183.

Boszormenyi:1993:CMO

[Bös93] László Böszörményi. A comparison of modula-3 and oberon-2. Structured Programming, 14(1): 15–22, 1993. CODEN STPGEM. ISSN 0935-1183.

Boszormenyi:1998:WJM

[Bös98] László Böszörményi. Why Java is not my favorite first-course language. Software — Concepts and Tools, 19(3):141–145, March 1998. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic). URL https:/

                                        /link.springer.com/article/
                                        10.1007/s003780050017.

Blaschek:1989:COO

[BPS89] Günther Blaschek, Gustav Pomberger, and Alois Stritzinger. A comparison of object-oriented programming languages. Structured Programming, 10(4):187–197, 1989. CODEN STPGEM. ISSN 0935-1183.

Braunl:1989:SSP

[Brä89] Thomas Bräunl. Structured SIMD programming in parallaxis. Structured Programming, 10(3):121–132, 1989. CODEN STPGEM. ISSN 0935-1183.

Bredereke:1996:DFI

[Bre96] Jan Bredereke. Detection of feature interactions in intelligent networks by verification. Software — Concepts and Tools, 17(3):121–139, 1996. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

BrinchHansen:1990:LSR

[Bri90] Per Brinch Hansen. The linear search rediscovered. Structured Programming, 11(1):53– 55, 1990. CODEN STPGEM. ISSN 0935-1183. URL http:

//brinch-hansen.net/papers/

1990a.pdf. See [DF89].

Bron:1990:SGA

[Bro90] Coenraad Bron. Symmetric (and generic) algorithms for height balanced trees. Structured Programming, 11(2):59–68, 1990. CODEN STPGEM. ISSN 0935-1183.

Bron:1992:LAA

[Bro92] Coenraad Bron. Letter to the authors of the article ’on intermediate variables’. Structured Programming, 13(1):51–53, 1992. CODEN STPGEM. ISSN 0935-

Broy:1997:TMC

[Bro97] Manfred Broy. Towards a mathematical concept of a component and its use. Software — Concepts and Tools, 18(3):137– ??, 1997. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Broy:1998:UMC

[Bro98] Manfred Broy. A uniform mathematical concept of a component. Software — Concepts and Tools, 19(1):57–59, March 1998. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic). URL https://

link.springer.com/article/
10.1007/s003780050008.

Back:1991:SRA

[BS91] Ralph-Johan Back and Kaisa Sere. Stepwise refinement of action systems. Structured Programming, 12(1):17–30, 1991. CODEN STPGEM. ISSN 0935-1183.

Back:1996:ASM

[BS96] Ralph-Johan Back and Kaisa Sere. From action systems to modular systems. Software — Concepts and Tools, 17(1):26– 39, 1996. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Borger:2000:IPJ

[BS00] Egon Börger and Wolfram Schulte. Initialization problems for Java. Software — Concepts and Tools, 19(4):175–178, October 2000. CODEN SCOTE5.

ISSN 0945-8115 (print), 1432-2188 (electronic). URL https:/

/link.springer.com/article/
10.1007/s003789900003.

Backhouse:1997:MPS

[BVW97] Roland Carl Backhouse, Richard Verhoeven, and Olaf Weber. Math/pad: a system for online preparation of mathematical documents. Software — Concepts and Tools, 18(2):80– ??, 1997. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Cregut:1997:POO

[CC97] Xavier Crégut and Bernard Coulette. PBOOL: an objectoriented language for definition and reuse of enactable processes. Software — Concepts and Tools, 18(2):47–62, 1997. CO-DEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Celko:1989:OAS

[Cel89] Joe Celko. Observations about student programming practices. Structured Programming, 10 (4):215–??, 1989. CODEN STPGEM. ISSN 0935-1183.

Cramer:1991:TFB

[CFGGR91] Joachim Cramer, Werner Fey, Michael Goedicke, and Martin Große-Rhode. Towards a formally based component description language — a foundation for reuse. Structured Programming, 12(2):91–110, 1991. CO-DEN STPGEM. ISSN 0935-

Champlain:1991:SOB

[Cha91] Michel De Champlain. Synapse: an object-based real-time programming language. Structured Programming, 12(3):145– 156, 1991. CODEN STPGEM. ISSN 0935-1183.

Christiansen:1989:SSI

[Chr89] Henning Christiansen. Structure sharing in incremental systems. Structured Programming, 10(4):169–186, 1989. CODEN STPGEM. ISSN 0935-1183.

Christensen:1993:SPL

[Chr93] H. Christensen. Synthesis of programs from logic specifications using programming methodology. Structured Programming, 14(4):173–186, 1993. CODEN STPGEM. ISSN 0935-1183.

Cowan:1993:ADV

[CIdLS93] Donald D. Cowan, Roberto Ierusalimschy, Carlos José Pereira de Lucena, and Terry M. Stepien. Abstract data views. Structured Programming, 14(1):1–14, 1993. CODEN STPGEM. ISSN 0935-1183.

Chen:1997:EDC

[CL97] Tsong Yueh Chen and Chi Keen Low. Error detection in C++ through dynamic data flow analysis. Software — Concepts and Tools, 18(1):1–13, 1997. CO-DEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Ciampolini:1996:OID

[CLM96] Anna Ciampolini, Evelina Lamma, and Paola Mello. An optimized implementation of a dynamic modular logic language. Software — Concepts and Tools, 17(4):148–162, 1996. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Corradi:1997:OOA

[CLZ97] Antonio Corradi, Letizia Leonardi, and Franco Zambonelli. An object-oriented approach to structured parallel programming. Software — Concepts and Tools, 18(3):110–124, 1997. CO-DEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Celi:1992:GMV

[CM92] Maria Luisa Celi and Vitaliano Milanese. Graphics modeling and visualization: an amalgamating approach using OBERON. Structured Programming, 13(4):179–192, 1992. CO-DEN STPGEM. ISSN 0935-1183.

Cook:1995:SAA

[CM95] Robert P. Cook and Richard G.

McDaniel. The StarLite algorithm animator. Software — Concepts and Tools, 16(1):1– 11, 1995. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Cleaveland:1996:MVD

[CNSL96] Rance Cleaveland, V. Natarajan, Steve Sims, and Gerald Lüttgen. Modeling and verifying distributed systems using priorities: a case study. Software — Concepts and Tools, 17(2):

50–62, 1996. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Cook:1996:TEA

[Coo96] Robert P. Cook. The type extensible architecture of a simple database system. Software — Concepts and Tools, 17(4):141– 147, 1996. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Cook:1998:ALC

[Coo98] Robert P. Cook. Anon-line C++ class: implementation and experience. Software — Concepts and Tools, 19(2):61–69, March 1998. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic). URL https:/

                                        /link.springer.com/article/
                                        10.1007/s003780050009.

Cook:1996:IOT

[CP96] Robert P. Cook and Krishnan Pillaipakkamnatt. An introduction to opaque types. Software — Concepts and Tools, 17(1): 1–5, 1996. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Cao:1997:SES

[CP97] Jiannong Cao and Mathew Pole. A software environment for simulating distributed taskscheduling algorithms. Software — Concepts and Tools, 18(3):125–136, 1997. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Cotronis:1994:LSC

[CRE94] John Yiannis Cotronis, Apostolos Nikolaos Refenes, and Eugene Eberbach. Language support for concurrent symbolic and numeric systems. Software — Concepts and Tools, 15(3):105– 124, 1994. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Crelier:1995:EMI

[Cre95] Régis Crelier. Extending module interfaces without invalidating clients. Software — Concepts and Tools, 16(2):49–62, 1995. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Childs:1997:ALP

[CS97] Bart Childs and Johannes Sametinger. Analysis of literate programs from the viewpoint of reuse. Software — Concepts and Tools, 18(1):35–??, 1997. CO-DEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Capel:1994:OBT

[CT94] Manuel I. Capel and José M.

Troya. An object-based tool and methodological approach for distributed programming. Software — Concepts and Tools, 15(4): 177–??, 1994. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Deinhammer:1997:NAC

[Dei97] Guido Deinhammer. A new adaptive coding scheme and its application to LZ-based data compression. Software — Concepts and Tools, 18(4):157–168, 1997. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Deimel:1998:SBF

[Dei98] Anton Deimel. The SAPR/3 business framework. Software — Concepts and Tools, 19(1):29–36, March 1998. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic). URL https:/

                                        /link.springer.com/article/
                                        10.1007/s003780050005.

Dijkstra:1989:LSR

[DF89] Edsger W. Dijkstra and W. H. J. Feijen. The linear search revisited. Structured Programming, 10(1):5–9, 1989. CODEN STPGEM. ISSN 0935-1183.

Doberkat:1992:IPS

[Dob92a] Ernst-Erich Doberkat. Integrating persistence into a set-oriented prototyping language. Structured Programming, 13(3):137– 153, 1992. CODEN STPGEM. ISSN 0935-1183.

Dobler:1992:APC

[Dob92b] Heinz Dobler. Ada on personal computers — some experiences. Structured Programming, 13(4):193–202, 1992. CODEN STPGEM. ISSN 0935-1183.

Doberkat:1996:LSH

[Dob96] Ernst-Erich Doberkat. A language for specifying hyperdocuments. Software — Concepts and Tools, 17(4):163–172, 1996. CO-DEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Dolinsky:1998:ACL

[DP98] Jens-Uwe Dolinsky and Thorsten Pawletta. Alightweight class library for extended persistent object management in C++. Software — Concepts and Tools, 19(2):71–79, March 1998. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic). URL https://

link.springer.com/article/
10.1007/s003780050010.

Egorov:1993:YAA

[EBP93] Igor I. Egorov, Ruslan P. Bogatyrev, and Dmitry L. Petrovichev. Yet another approach to modula-2 implementation of exception handling mechanism. Structured Programming, 14(1): 23–36, 1993. CODEN STPGEM. ISSN 0935-1183.

Ejiogu:1990:BSP

[Eji90] Lem O. Ejiogu. Beyond structured programming: an introduction to the principles of applied software metrics. Structured Programming, 11(1):27–44, 1990. CODEN STPGEM. ISSN 0935-1183.

Eppler:1993:STL

[Epp93] Reinhard Eppler. Standard type LIST for imperative languages. Structured Programming, 14(2): 56–61, 1993. CODEN STPGEM. ISSN 0935-1183.

Ephraim:1993:ASC

[ETJJ93] M. J. Ephraïm, Ad H. M. Thiers, A. Janner, and T. Janssen. Applying Smalltalk-80 and C++ to crystal symmetry analysis. Structured Programming, 14(3):119–135, 1993. CODEN STPGEM. ISSN 0935-1183.

Franz:1993:CUS

[Fra93] Michael Franz. The case for universal symbol files. Structured Programming, 14(3):136– 147, 1993. CODEN STPGEM. ISSN 0935-1183.

Franz:1995:PET

[Fra95] Michael Franz. Protocol extension: a technique for structuring large extensible softwaresystems. Software — Concepts and Tools, 16(2):86–94, 1995. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Franz:1997:PLL

[Fra97] Michael Franz. The programming language Lagoona — a fresh look at object-orientation. Software — Concepts and Tools, 18(1):14–26, 1997. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Fuchs:1992:HLE

[Fuc92] Norbert E. Fuchs. Hoare logic, executable specifications, and logic programs. Structured Programming, 13(3):129–135, 1992. CODEN STPGEM. ISSN 0935-1183.

Fuchs:1993:IAD

[Fuc93] Norbert E. Fuchs. An improved algorithm for deriving the decomposition tree of a flowgraph. Structured Programming, 14(3):93–101, 1993. CODEN STPGEM. ISSN 0935-1183.

Grechenig:1993:CIO

[GB93] Thomas Grechenig and Stefan Biffl. The challenge of introducing the object-oriented paradigm — an empirical investigation of a software-engineering course. Structured Programming, 14(4):187–198, 1993. CODEN STPGEM. ISSN 0935-1183.

Gomard:1991:CGP

[GJ91] Carsten K. Gomard and Neil D.

Jones. Compiler generation by partial evaluation: a case study. Structured Programming, 12(3):123–144, 1991. CODEN STPGEM. ISSN 0935-1183.

Graham:1992:DIC

[GK92] T. C. Nicholas Graham and Gerd Kock. Domesticating imperative constructs for a functional world. Structured Programming, 13(2): 75–88, 1992. CODEN STPGEM. ISSN 0935-1183.

Giordano:1994:SPL

[GM94] Laura Giordano and Alberto Martelli. Structured prolog: a language for structured logic programming. Software — Concepts and Tools, 15(3):125–??, 1994. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Goldberg:1998:RBM

[Gol98] Adele Goldberg. A reuse business model. Software — Concepts and Tools, 19(1):11–13, March 1998. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic). URL https:/

                                        /link.springer.com/article/
                                        10.1007/s003780050003.

Goos:1997:SKL

[Goo97] Gerhard Goos. Sather-K — the language. Software — Concepts and Tools, 18(3):91–109, 1997. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Gaskell:1995:SAE

[GP95a] Craig Gaskell and Roger Phillips. Software architecture of the executable graphical specification tool EGS1. Software — Concepts and Tools, 16(3):124–135, 1995. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Gravell:1995:FMO

[GP95b] Andrew M. Gravell and Chris H. Pratten. Formal methods and open systems. Software — Concepts and Tools, 16(4):183– ??, 1995. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Grosch:1990:ECE

[Gro90] Josef Grosch. Efficient and comfortable error recovery in recursive descent parsers. Structured Programming, 11(3):129– 140, 1990. CODEN STPGEM. ISSN 0935-1183.

Grogono:1991:IDO

[Gro91a] Peter Grogono. Issues in the design of an object oriented programming language. Structured Programming, 12(1):1–15, 1991. CODEN STPGEM. ISSN 0935-1183.

Grosch:1991:TSD

[Gro91b] Josef Grosch. Tool support for data structures. Structured Programming, 12(1):31–38, 1991. CODEN STPGEM. ISSN 0935-1183.

Gupta:1990:DCR

[Gup90] Rajiv Gupta. Debugging code reorganized by a trace scheduling compiler. Structured Programming, 11(3):141–150, 1990. CODEN STPGEM. ISSN 0935-1183.

Gutknecht:1989:VRM

[Gut89] Jürg Gutknecht. Variations on the role of module interfaces. Structured Programming, 10(1): 40–46, 1989. CODEN STPGEM. ISSN 0935-1183.

Gutknecht:1993:TPL

[Gut93] Jürg Gutknecht. The tragedy of programming language development. Structured Programming, 14(2):49–55, 1993. CO-DEN STPGEM. ISSN 0935-1183.

Gutknecht:1994:OSV

[Gut94] Jürg Gutknecht. Oberon System 3: Vision of a future software technology. Software — Concepts and Tools, 15(1):26– 33, 1994. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Gries:1990:TNL

[GV90] David Gries and Dennis M. Volpano. The transform — a new language construct. Structured Programming, 11(1):1–10, 1990. CODEN STPGEM. ISSN 0935-1183.

Hartel:1990:CTG

[Har90] Pieter H. Hartel. A comparison of three garbage collection algorithms. Structured Programming, 11(3):117–128, 1990. CODEN STPGEM. ISSN 0935-1183.

Henn:1998:ISF

[Hen98] Juergen Henn. IBM San Francisco. Software — Concepts and Tools, 19(1):37–48, March 1998. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic). URL https://

link.springer.com/article/
10.1007/s003780050006.

Heuberger:1998:MUI

[Heu98] Philipp Heuberger. The minimal user interface of a simple refinement tool. Software — Concepts and Tools, 19(2):89–95, March 1998. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic). URL https:/

/link.springer.com/article/
10.1007/s003780050012.

Hirakawa:1994:VLS

[HI94] Masahito Hirakawa and Tadao Ichikawa. Visual language studies — a perspective. Software — Concepts and Tools, 15(2): 61–67, 1994. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Hull:2000:TVA

[HNR00] M. Elizabeth C. Hull, Peter N.

Nicholl, and Niall Rooney. Towards a visual approach for component-based software development. Software — Concepts and Tools, 19(4):154–160, October 2000. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic). URL https:/

/link.springer.com/article/
10.1007/s003789900001.

Holzmann:1996:EFD

[Hol96] Gerard J. Holzmann. Early fault detection tools. Software — Concepts and Tools, 17(2):63– 69, 1996. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Heeb:1991:IVL

[HP91] Beat Heeb and Cuno Pfister. On intermediate variables and local procedures as parameters. Structured Programming, 12(1):39–42, 1991. CODEN STPGEM. ISSN 0935-1183.

Hrebabetzky:1997:CLC

[Hre97] Fran Hrebabetzky. Closed-loop control with oberon. Software — Concepts and Tools, 18(2): 73–79, 1997. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Hougaard:1995:TIT

[HSA95] Ole I. Hougaard, Michael I. Schwartzbach, and Hosein Askari. Type inference for turbo pascal. Software — Concepts and Tools, 16(4):160–169, 1995. CO-DEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Klein:1992:HPP

[KK92] Eduard Klein and Kai Koskimies. How to pipeline parsing with parallel semantic analysis. Structured Programming, 13(3):99– 107, 1992. CODEN STPGEM. ISSN 0935-1183.

Knuth:1994:MIL

[Knu94] Donald E. Knuth. Mini-indexes for literate programs. Software — Concepts and Tools, 15(1): 2–11, 1994. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Kofler:1993:RI

[Kof93] Thomas Kofler. Robust iterators in ET++. Structured Programming, 14(2):62–85, 1993. CO-DEN STPGEM. ISSN 0935-1183.

Kimura:1994:GDE

[KSA94] Takayuki Dan Kimura, Samudra Sengupta, and Ajay Apte. A graphic diagram editor for pen computers. Software — Concepts and Tools, 15(2):82–??, 1994. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Keller:1995:ESB

[KSL^{+}95] Rudolf K. Keller, Xijin Shen, Richard Lajoie, Marianne Ozkan, and Tao Tao. Environment support for business reengineering: The macrotec approach. Software — Concepts and Tools, 16(1):31–40, 1995. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Klaeren:1990:CMI

[KT90] Herbert Klaeren and Peter Thiemann. A clean modula-2 interface to abstract data types. Structured Programming, 11(2): 69–78, 1990. CODEN STPGEM. ISSN 0935-1183.

Lampe:1997:SAM

[Lam97] Jürgen Lampe. Statistics about modules of the oberon system. Software — Concepts and Tools, 18(1):27–34, 1997. CO-DEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Lampe:1998:DGD

[Lam98] Jürgen Lampe. Depot4 — a generator for dynamically extensible translators. Software — Concepts and Tools, 19(2):97–108, March 1998. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic). URL https:/

/link.springer.com/article/
10.1007/s003780050013.

Lee:2000:MBR

[Lee00]  M. H. Lee.  Model-based

reasoning: a principled approach for software engineering. Software — Concepts and Tools, 19(4):179–189, October 2000. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic). URL https:/

                                        /link.springer.com/article/
                                        10.1007/s003780000005.

Lins:1989:FLA

[Lin89a] Charles Lins. A first look at literate programming. Structured Programming, 10(1):60–62, 1989. CODEN STPGEM. ISSN 0935-1183.

Lins:1989:ILP

[Lin89b] Charles Lins. An introduction to literate programming. Structured Programming, 10(2):107– 112, 1989. CODEN STPGEM. ISSN 0935-1183.

Lowe:1996:BFN

[Low96] Gavin Lowe. Breaking and fixing the needham-schroeder publickey protocol using FDR. Software — Concepts and Tools, 17(3):93–102, 1996. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Lokpo:1992:ROL

[LP92] Bamba Ibrahim Lokpo and Gérard Padiou. Reusability in the occam language. Structured Programming, 13(2):65–74, 1992. CODEN STPGEM. ISSN 0935-1183.

Lalis:1995:POO

[LS95] Spyros Lalis and Beverly A. Sanders. From persistent objects to object transmission in distributed systems using oberon. Software — Concepts and Tools, 16(1):12–19, 1995. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Marais:1991:GUI

[Mar91] Johannes L. Marais. The GAD-GETS user interface management system. Structured Programming, 12(2):75–90, 1991. CODEN STPGEM. ISSN 0935-1183.

McCartney:1997:USE

[McC97] Paul McCartney. A usability study of end-user construction of distributed multimedia applications. Software — Concepts and Tools, 18(2):63–72, 1997. CO-DEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Meyer:1989:SPO

[Mey89] Bertrand Meyer. From structured programming to objectoriented design: The road to Eiffel. Structured Programming, 10(1):19–39, 1989. CODEN STPGEM. ISSN 0935-1183.

Michelis:1993:RC

[MG93] Giorgio De Michelis and Maria Antonietta Grasso. Routines and conversations. Structured Programming, 14(3):110–118, 1993. CODEN STPGEM. ISSN 0935-1183.

McCartney:1995:EEU

[MGS95] Paul McCartney, Kenneth J.

Goldman, and David E. Saff. EUPHORIA: End-user construction of direct manipulation user interfaces for distributed applications. Software — Concepts and Tools, 16(4):147–159, 1995. CO-DEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Milanese:1991:AGL

[Mil91] Vitaliano Milanese. Aspects of a graphic language: a critical evaluation. Structured Programming, 12(3):157–??, 1991. CO-DEN STPGEM. ISSN 0935-1183.

Mansharamani:2000:RTG

[MKVM00] Rajesh Mansharamani, Prasad Kallepalli, Harsha Veerabhadraiah, and Benny Mathew. RV-GEN: a tool for generation of random variates. Software — Concepts and Tools, 19(4):161–167, October 2000. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic). URL https://

link.springer.com/article/
10.1007/s003789900002.

Morris:1990:PER

[Mor90] Joseph M. Morris. Programming by expression refinement: a sequence of examples. Structured Programming, 11(4):189– 198, 1990. CODEN STPGEM. ISSN 0935-1183.

Moser:1995:MOO

[Mos95]  Simon Moser.  Metamodels

for object oriented systems: a proposition of metamodels describing object oriented systems at consecutive levels of abstraction. Software — Concepts and Tools, 16(2):63–80, 1995. CO-DEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Miller:1990:SPP

[MS90] Russ Miller and Quentin F.

Stout. Seymour: a portable parallel programming language. Structured Programming, 11 (4):157–172, 1990. CODEN STPGEM. ISSN 0935-1183.

Mendelzon:1995:REV

[MS95] Alberto O. Mendelzon and Johannes Sametinger. Reverse engineering by visualizing and querying. Software — Concepts and Tools, 16(4):170–182, 1995. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Margaria:1996:F

[MS96] Tiziana Margaria and Bernhard Steffen. Foreword. Software — Concepts and Tools, 17(2): 49–??, 1996. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Mossenbock:1989:OOM

[MT89] Hanspeter Mössenböck and Josef Templ. Object Oberon — a modest object-oriented language. Structured Programming, 10(4):199–207, 1989. CODEN STPGEM. ISSN 0935-1183.

Mossenbock:1991:PLO

[MW91a] H. Mössenböck and N. Wirth.

The programming language Oberon-2. Structured Programming, 12(4):179–195, ???? 1991. CODEN STPGEM. ISSN 0935-

Mossenbock:1991:DBO

[MW91b] Hanspeter Mössenböck and Niklaus Wirth. Differences between Oberon and Oberon-2. Structured Programming, 12 (4):175–178, 1991. CODEN STPGEM. ISSN 0935-1183.

Navrat:1996:KCV

[NB96] Pavol Návrat and Mária Bieliková. Knowledge-controlled version selection in software configuration management. Software — Concepts and Tools, 17(1):40– ??, 1996. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Odersky:1989:MNA

[Ode89] Martin Odersky. MINOS: a new approach to the design of an input/output library for Modula-2. Structured Programming, 10(2):89–105, 1989. CODEN STPGEM. ISSN 0935-1183.

Padiou:1990:NCE

[Pad90] Gérard Padiou. Nested coroutines for exception handling in modula-2. Structured Programming, 11(2):79–84, 1990. CO-DEN STPGEM. ISSN 0935-1183.

Pomberger:1991:POS

[PBK^{+}91] Gustav Pomberger, Walter R. Bischofberger, Dieter Kolb, Wolfgang Pree, and Holger Schlemm. Prototyping-oriented software development — concepts and tools. Structured Programming, 12(1):43–60, 1991.

CODEN STPGEM. ISSN 0935-1183.

Parr:1998:DMP

[PC98] Gerard Parr and Kevin Curran. Dynamic multimedia protocol stacks. Software — Concepts and Tools, 19(2):81–88, March 1998. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic). URL https:/

/link.springer.com/article/
10.1007/s003780050011.

Pirklbauer:1992:SPM

[Pir92] Klaus Pirklbauer. A stuudyy of pattern-matching algorithms. Structured Programming, 13(2): 89–??, 1992. CODEN STPGEM. ISSN 0935-1183.

Pomberger:1989:EW

[Pom89] Gustav Pomberger. Editorial welcome. Structured Programming, 10(1):3–4, 1989. CODEN STPGEM. ISSN 0935-1183.

Pomberger:1990:E

[Pom90]  Gustav Pomberger.  Edito-

rial. Structured Programming, 11(4):155–156, 1990. CODEN STPGEM. ISSN 0935-1183.

Pomberger:1998:E

[Pom98]  Gustav Pomberger.  Edito-

rial. Software — Concepts and Tools, 19(1):1, March 1998. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic). URL https://

link.springer.com/article/
10.1007/s003780050001.

Pomberger:2000:DRS

[Pom00]  Gustav Pomberger.  Dear

readers, subscribers and authors. Software — Concepts and Tools, 19(4):153, October 2000. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic). URL https:/

/link.springer.com/article/
10.1007/s003780000007.

Pree:1995:AGF

[PPSS95] Wolfgang Pree, Gustav Pomberger, Albert Schappert, and Peter Sommerlad. Active guidance of framework development. Software — Concepts and Tools, 16(3):136–??, 1995. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Pomberger:1994:E

[PR94] Gustav Pomberger and Gerhard Rossbach. Editorial. Software — Concepts and Tools, 15(1): 1–??, 1994. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Pree:1997:CBS

[Pre97] Wolfgang Pree. Componentbased software development — a new paradigm in software engineering? Software — Concepts and Tools, 18(4):169–174, 1997. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Pritchard:1989:DOS

[Pri89]  Paul Pritchard.  A display-

oriented syntax notation. Structured Programming, 10(3):165– 168, 1989. CODEN STPGEM. ISSN 0935-1183.

Plasil:1998:AVD

[PS98] František Plášil and Michael Stal. An architectural view of distributed objects and components in CORBA, Java RMI and COM/DCOM. Software — Concepts and Tools, 19(1):14–28, March 1998. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic). URL https:/

/link.springer.com/article/
10.1007/s003780050004.

Radenski:1998:ME

[Rad98] Atanas Radenski. Module embedding. Software — Concepts and Tools, 19(3):122–129, March 1998. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic). URL https:/

/link.springer.com/article/
10.1007/s003780050015.

Rossello:1989:DSE

[RBP89] Celestí Rosselló, José L. Balcázar, and Ricardo Pena. Deriving specifications of embeddings in recursive program design. Structured Programming, 10(3):133– 145, 1989. CODEN STPGEM. ISSN 0935-1183.

Rechenberg:1990:PLT

[Rec90] Peter Rechenberg. Programming languages as thought models. Structured Programming, 11(3):105–116, 1990. CODEN STPGEM. ISSN 0935-1183.

Reichenberger:1994:CTS

[Rei94] Christoph Reichenberger. Concepts and techniques for software version control. Software — Concepts and Tools, 15(3):97– 104, 1994. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Revesz:1991:TOM

[Rév91] György E. Révész. On translating ordinary mathematical notation. Structured Programming, 12(3):115–122, 1991. CODEN STPGEM. ISSN 0935-1183.

Rizman:1993:BMS

[Riz93]  Krista Rizman.  Building

more structured, understandable, reusable and changeable object-oriented software. Structured Programming, 14(3):102– 109, 1993. CODEN STPGEM. ISSN 0935-1183.

Reinfelds:1990:DPF

[RM90] Juris Reinfelds and Richard Miller. The development of a program to find the length of the longest m-sequence in a string. Structured Programming, 11(1): 45–52, 1990. CODEN STPGEM. ISSN 0935-1183.

Rose:1992:NRA

[Ros92] Julian Rose. A new rigorous approach for modelling and refining concurrent behaviour in JSD specifications. Structured Programming, 13(1):11–22, 1992. CODEN STPGEM. ISSN 0935-1183.

Sakkinen:1992:DSC

[Sak92] Markku Sakkinen. The darker side of C++ revisited. Structured Programming, 13(4):155– 178, 1992. CODEN STPGEM. ISSN 0935-1183.

Salomon:1997:IPG

[Sal97] Ralf Salomon. Improving the performance of genetic algorithms through derandomization. Software — Concepts and Tools, 18(4):175–??, 1997. CO-DEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Sanders:2000:UAA

[San00]  Beverly A. Sanders.  Using

atomic await commands to develop concurrent programs in Java. Software — Concepts and Tools, 19(4):168–174, October 2000. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic). URL https:/

/link.springer.com/article/
10.1007/s003789900004.

Schaffer:1990:CHO

[Sch90] Bruno Schäffer. Chiron — a homogeneous object-oriented language. Structured Programming, 11(4):173–188, 1990. CODEN STPGEM. ISSN 0935-1183.

Schnorf:1991:DRS

[Sch91] Peter Schnorf. Design of a reusable symbol table abstraction. Structured Programming, 12(2):63–74, 1991. CODEN STPGEM. ISSN 0935-1183.

Schmitz:1992:UIE

[Sch92] Lothar Schmitz. Using inheritance to explore a family of algorithms. Structured Programming, 13(2):55–64, 1992. CO-DEN STPGEM. ISSN 0935-1183.

Suda:1991:TSI

[SD91] Peter Suda and Hans Delfs. TOROS — a system integrating object and rule-oriented programming in a C/UNIX environment. Structured Programming, 12(4):207–216, 1991. CODEN STPGEM. ISSN 0935-1183.

Sikora:1993:OOI

[SGHR93] Herman Sikora, Paul Grünbacher, Günter Hummel, and Siegfried Reich. Object-oriented implementation apsects of an abstract relational database interface. Structured Programming, 14(2):86–??, 1993. CODEN STPGEM. ISSN 0935-1183.

Sajaniemi:1998:VSP

[SI98] Jorma Sajaniemi and Kari Ikonen. VinEd — a system for program manipulation through user-definable simultaneous views. Software — Concepts and Tools, 19(3):130–140, March 1998. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic). URL https:/

/link.springer.com/article/
10.1007/s003780050016.

Steffen:1996:HAV

[SMC96] Bernhard Steffen, Tiziana Margaria, and Andreas Claßen. Heterogeneous analysis and verification for distributed systems. Software — Concepts and Tools, 17(1):13–25, 1996. CO-DEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Steffen:1996:IFK

[SMCB96] Bernhard Steffen, Tiziana Margaria, Andreas Claßen, and Volker Braun. Incremental formalization: a key to industrial success. Software — Concepts and Tools, 17(2):78–??, 1996. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Spinellis:1993:IHL

[Spi93] Diomidis Spinellis. Implementing Haskell: Language implementation as a tool building exercise. Structured Programming, 14(1):37–48, 1993. CO-DEN STPGEM. ISSN 0935-

1183. URL http://kerkis.
math.aegean.gr/~dspin/pubs/
jrnl/1993-StrProg-Haskell/
html/exp.html.

Smith:1992:MCS

[SS92] Lisa M. Smith and Mansur H.

Samadzadeh. Measuring complexity and stability of WEB programs. Structured Programming, 13(1):35–50, 1992. CO-DEN STPGEM. ISSN 0935-1183.

Stary:1996:IWR

[Sta96] Christian Stary. Integrating workflow representations into user interaface design representations. Software — Concepts and Tools, 17(4):173–??, 1996. CO-DEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Spoelder:1993:CCC

[SUL93] Hans J. W. Spoelder, A. H. Ullings, and C. N. Liu. Coupling a C++ class with existing software: a case study with BLAS. Structured Programming, 14(4):150–157, 1993. CODEN STPGEM. ISSN 0935-1183.

Szyperski:1998:ECS

[Szy98]  Clemens Szyperski.  Emerg-

ing component software technologies — a strategic comparison. Software — Concepts and Tools, 19(1):2–10, March 1998. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic). URL https://

                                        link.springer.com/article/
                                        10.1007/s003780050002.

Templ:1991:DIS

[Tem91] Josef Templ. Design and implementation of SPARC-Oberon. Structured Programming, 12 (4):197–206, 1991. CODEN STPGEM. ISSN 0935-1183.

Turner:1995:SBA

[TR95] Christopher David Turner and David J. Robson. A state-based approach to the testing of classbased programs. Software — Concepts and Tools, 16(3):106– 112, 1995. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Tretmans:1996:TGI

[Tre96] Jan Tretmans. Test generation with inputs, outputs and repetitive quiescence. Software — Concepts and Tools, 17(3):103– 120, 1996. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Tung:1989:SMLb

[Tun89] Sho-Huan Tung. A structured method for literate programming. Structured Programming, 10(2):113–120, 1989. CODEN STPGEM. ISSN 0935-1183.

Turski:1992:ECP

[Tur92] Wladyslaw M. Turski. Extending the computing paradigm. Structured Programming, 13(1):1–10, 1992. CODEN STPGEM. ISSN 0935-1183.

Toleman:1998:SED

[TW98] Mark A. Toleman and Jim Welsh. Systematic evaluation of design choices for software development tools. Software — Concepts and Tools, 19(3):109–121, March 1998. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic). URL https:/

/link.springer.com/article/
10.1007/s003780050014.

Volpano:1995:TCC

[Vol95] Dennis M. Volpano. Type checking coherent overloading. Software — Concepts and Tools, 16(2):81–85, 1995. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Ward:1994:LOP

[War94] Martin P. Ward. Languageoriented programming. Software — Concepts and Tools, 15(4):147–161, 1994. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Wildman:2000:CAD

[WFC00] Luke Wildman, Colin Fidge, and David Carrington. Computeraided development of a real-time program. Software — Concepts and Tools, 19(4):190–202, October 2000. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic). URL https:/

                                        /link.springer.com/article/
                                        10.1007/s003780000006.

Weinand:1989:DIS

[WGM89] André Weinand, Erich Gamma, and Rudolph Marty. Design and implementation of ET++, a seamless object-oriented application framework. Structured Programming, 10(2):63–87, 1989. CODEN STPGEM. ISSN 0935-1183.

Welsh:1994:SDC

[WH94] Jim Welsh and Jun Han. Software documents: Concepts and tools. Software — Concepts and Tools, 15(1):12–25, 1994. CO-DEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Wirth:1989:DSS

[Wir89] Niklaus Wirth. Designing a system from scratch. Structured Programming, 10(1):10–18, 1989.

CODEN STPGEM. ISSN 0935-1183.

Wirth:1996:TVT

[Wir96] Niklaus Wirth. Tasks versus threads: an alternative multiprocessing paradigm. Software — Concepts and Tools, 17(1): 6–12, 1996. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Waeselynck:1994:SDA

[WP94] Hélène Waeselynck and Dietmar Pfahl. System dynamics applied to the modelling of software projects. Software — Concepts and Tools, 15(4):162–176, 1994. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Welsh:1994:IST

[WY94] Jim Welsh and Yun Yang. Integration of semantic tools into document editors. Software — Concepts and Tools, 15(2):68– 81, 1994. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Xue:1997:DPK

[XD97] Jinyun Xue and Ruth Davis. A derivation and proof of Knuth’s binary to decimal conversion program. Software — Concepts and Tools, 18(4):149–156, 1997. CO-DEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Yeager:1989:HSA

[Yea89] Dorian P. Yeager. A historysensitive approach to the description of binary relations. Structured Programming, 10(3):157– 163, 1989. CODEN STPGEM. ISSN 0935-1183.

Yuen:1995:BPA

[YF95a] Chung-Kwong Yuen and Ming-Dong Feng. BaLinda plus: Adding objects to parallel languages. Software — Concepts and Tools, 16(3):95–105, 1995. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Yuen:1995:ICS

[YF95b] Chung-Kwong Yuen and Ming-Dong Feng. Iterative computation and speculative processing. Software — Concepts and Tools, 16(1):41–??, 1995. CO-DEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Zimmer:1995:TSC

[Zim95] J. Adrian Zimmer. Type substitution in compiled code. Software — Concepts and Tools, 16(1): 20–30, 1995. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic).

Zhou:1998:NNC

[ZWC98] Yan Zhou, Dawn Wilkins, and Robert P. Cook. Neural network control for a fire-fighting robot. Software — Concepts and Tools, 19(3):146–152, March 1998. CODEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic). URL https://

link.springer.com/article/
10.1007/s003780050018.